Simplify sin 2 y / (1 - cos y)

Simplify


sin2y / (1 - cos y)


 

Grade:11

1 Answers

Vijay Luxmi Askiitiansexpert
14 years ago

       plz check this : 

 
          (sin2 y)/(1-cos y)=(1-cos2 y)/(1-cos y)


        = (12 - cos2y)/(1-cos y)


        = (1 + cos y)


        = 1 + cos y


        =  ans.